TRIMMING

DARYL COOPER AND DARREN D. LONG

Abstract. We consider the question of deciding whether or not certain subgroups of certain cocompact Kleinian groups are geometrically finite or infinite. This is reduced to a question about similarity interval exchange maps. The concept of trimming produces an practical algorithm to settle the latterquestion. This algorithm has been implemented on a computer and the results are perplexing.